OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form.

OpenStudy (mr.math):

The slope-intercept form is \(y=mx+b\). In our case: \[y-4=-2(x-5) \implies y-4=-2x+10 \implies y=-2x+14.\]

OpenStudy (anonymous):

Please engage instead of just giving the answer.

OpenStudy (mr.math):

Whoops! Someone is getting mad!

OpenStudy (anonymous):

Wow, you get abusive because someone's trying to teach you?

OpenStudy (anonymous):

YUPPPPP

OpenStudy (anonymous):

Stay classy.

OpenStudy (mr.math):

I'm outta here before I get hit by erikasipe :P

OpenStudy (anonymous):

You too. [;

myininaya (myininaya):

Please don't be so mean and use inappropriate language, erikasipe. I think ktklown was just wanting to help.